Answer Question 3 please?(b) Use a = 2v5 and solve. (c) Use a = 6 and solve. 2. Solve the spring mass system y + y = u(t) + 6(t - A) y(0) = 0 Y(0= 0
Answer #1

The given function a(t) = { 0, for ostal 2t-2 for Kt 52 4-t for 22 AL 4 o for 468 we Sketch the fingetion gets below: weL CO

writting in terms of step finitions: get) = 0 for tt [0, 1) get) = 24-2 for A e [12] g(t) = 4-t for de [2, 4] gt) so for d E

Stest at a test -s est et =$(++)est to set-zest - 25 ? (43) St 72 =-2/)225 6+38] + Ees -- Le 250+5)-e 5,4-HEE ET +[* er* = f[..L{ ger} = -327(*)-&J + $(2-3):246-25 +38957 3 etso 3 eozes AS ut lat sens?